### v1

# BhasaAnuvaad: A Speech Translation Dataset for 14 Indian Languages

Abstract
Automatic Speech Translation (AST) datasets for Indian languages remain critically scarce, with public resources covering fewer than 10 of the 22 official languages. This scarcity has resulted in AST systems for Indian languages lagging far behind those available for high-resource languages like English. In this paper, we first evaluate the performance of widely-used AST systems on Indian languages, identifying notable performance gaps and challenges. Our findings show that while these systems perform adequately on read speech, they struggle significantly with spontaneous speech, including disfluencies like pauses and hesitations. Additionally, there is a striking absence of systems capable of accurately translating colloquial and informal language, a key aspect of everyday communication. To this end, we introduce BhasaAnuvaad, the largest publicly available dataset for AST involving 14 scheduled Indian languages spanning over 44,400 hours and 17M text segments. BhasaAnuvaad contains data for English speech to Indic text, as well as Indic speech to English text. This dataset comprises three key categories: (1) Curated datasets from existing resources, (2) Large-scale web mining, and (3) Synthetic data generation. By offering this diverse and expansive dataset, we aim to bridge the resource gap and promote advancements in AST for low-resource Indian languages, especially in handling spontaneous and informal speech patterns.
BhasaAnuvaad: A Speech Translation Dataset for 14 Indian Languages
Sparsh Jain1  Ashwin Sankar [^1] footnotemark: 1  Devilal Choudhary4thanks:   Equal Contribution.thanks:   Work done during internship at AI4Bharat.  Dhairya Suman311footnotemark: 122footnotemark: 2 Nikhil Narasimhan [^1] footnotemark: 1Mohammed Safi Ur Rahman Khan1,2 Anoop Kunchukuttan1,5Mitesh M Khapra1,2Raj Dabre1,2,6,7thanks:   Corresponding Author: raj.dabre@cse.iitm.ac.in 1 Nilekani Centre at AI4Bharat  2Indian Institute of Technology, Madras 3Indian Institute of Technology, Delhi  4Delhi Technological University 5Microsoft  6NICT, Japan  7Indian Institute of Technology, Bombay [Uncaptioned image] huggingface.co/BhasaAnuvaad   [Uncaptioned image] github.com/BhasaAnuvaad
1 Introduction
Figure 1: Per language statistics of the BhasaAnuvaad dataset for Speech Translation across 14 Indian languages, depicting number of thousands of hours paired with text translations.
Automatic Speech Translation (AST) has become crucial in enabling communication across languages and cultures by breaking down language barriers. Traditionally, speech translation systems have relied on cascaded architectures, where automatic speech recognition (ASR) is followed by machine translation (MT). However, recent advances have led to more integrated end-to-end (E2E) models Babu et al. (2021); Pratap et al. (2024); Radford et al. (2023); Communication et al. (2023) that directly translate speech from one language into text in another. Additionally, Audio-LLM-based systems have emerged Wu et al. (2023); Chu et al. (2023); Fathullah et al. (2023), further advancing the field. These improvements have mainly benefited high-resource languages like English and European languages, where abundant data and well-established benchmarks have driven significant progress.
Figure 2: Overview of the creation of BhasaAnuvaad.
In contrast, progress in AST for Indian languages has been much slower. India, with its vast linguistic diversity—22 officially recognized languages and numerous dialects—poses a complex challenge for building effective AST systems. Indian languages feature distinct phonetic systems Mujadia and Sharma (2023), frequent code-switching Shankar et al. (2024), and varied syntactic structures. Additionally, the limited availability of high-quality datasets for Indian languages, exacerbated by the limited representation of these languages on the web, hampers model development.
A critical issue in AST research is the predominant focus on evaluating systems using read speech benchmarks like FLUERS Conneau et al. (2022). While these benchmarks are useful, they fail to capture the complexities of spontaneous, real-world speech. Unlike read speech, spontaneous speech includes informal language, hesitations, extensive code-switching Shankar et al. (2024), and unpredictable structures, all of which are essential for building systems that can handle real-world use cases. Additionally in the case of FLEURS, the speech content has been recorded by only three native speaker, which doesn’t efficiently capture the dialectal variations present in the languages across India. Hence, systems that perform well on read speech often struggle with spontaneous speech, limiting their real-world applicability.
In this work, we first evaluate various state-of-the-art AST systems, across both Direct and Cascaded paradigms, on read speech (using FLUERS Conneau et al. (2022)) and a newly created benchmark, Indic-Spontaneous-Synth, designed to represent spontaneous speech in Indian languages. Our results show that while existing models perform well on read speech, they significantly underperform on spontaneous speech. This highlights the need for more comprehensive datasets, training objectives, and evaluations that better reflect real-world conditions.
To address this gap, we introduce BhasaAnuvaad, the largest Indic-language AST dataset spanning over 44,400 hours of speech and 17M text segments (Figure 1 and 3). This dataset includes mined parallel speech data, high-quality curated parallel speech data collected from minable corpora, and large-scale synthetic ST data generated using IndicVoices Javed et al. (2024), specifically designed to capture spontaneous speech in Indic languages (Figure 2). Out of the 44,400 hours, around 36,600 hours of data have been created through this project and released into the public domain. This includes about 7,400 hours of data from the SEAMLESS-Align project, where we replicated the intricate mining pipeline to collect the mined data. This diverse dataset provides the resources needed to develop more robust AST systems capable of handling spontaneous speech. In summary, our contributions are: (i) Indic-Spontaneous-Synth, a one of a kind created benchmark for evaluating spontaneous speech translations, (ii) an evaluation of popular AST baselines on both read and spontaneous benchmarks, and (iii) the release of BhasaAnuvaad, the largest Indic-language speech translation corpus, comprising over 44 thousand hours of data across 15 languages. All code and datasets developed as part of this work will be made publicly available to support future research in AST for Indic languages.
2 Related Works
Our work focuses on speech translation dataset creation for Indian languages. Cascaded AST. The traditional cascaded approach to Automatic Speech Translation (AST) has been a fundamental technique explored extensively in early research Stentiford and Steer (1988); Di Gangi et al. (2019a); Cheng et al. (2019); Bahar et al. (2020). This method connects Automatic Speech Recognition (ASR) and Machine Translation (MT) systems in sequence, which, while effective, is known to suffer from several limitations. Despite its straightforward and simple design, these systems are prone to issues such as cascaded error propagation Sperber and Paulik (2020); Beck et al. (2019), loss of prosodic information Bentivogli et al. (2021), and increased computational complexity Lee et al. (2022).
End-to-End AST. In contrast, end-to-end (E2E) models have gained traction due to their ability to be jointly trained, thereby eliminating the intermediate representations and reducing error propagation Weiss et al. (2017); Inaguma et al. (2021). However, these systems still face challenges, particularly with low-resource languages due to the lack of good quality data Sethiya and Maurya (2023). With the rise of modern Large Language Models (LLMs), recent works have started exploring the feasibility of using Audio-based LLMs (or Speech LLMs), which combine Speech Foundation models (SFMs) and LLMs Fathullah et al. (2023); Wu et al. (2023); Huang et al. (2023).
Indian Languages AST. While these advances have been significant for English and other languages, the lack of sufficient parallel datasets has hindered the progress in Indian Languages Mujadia and Sharma (2023); Mhaskar et al. (2023). Multilingual speech translation models like mSLAM Bapna et al. (2022), Whisper Radford et al. (2023), SeamlessM4T Communication et al. (2023) have made strides in addressing this gap, but their effectiveness in low-resource and real-world spontaneous speech settings remains largely unexplored. Although there are well-established benchmarks to evaluate text-based MT systems for Indian Languages Gala et al. (2023); Goyal et al. (2022), speech-based benchmarks are still underdeveloped Sethiya and Maurya (2023), with FLUERS Conneau et al. (2022) being the primary option. This gap, which we partially in this paper, is even more pronounced in the evaluation of spontaneous speech translation, where no benchmarks currently exist. This also makes it challenging to host shared tasks on AST for Indian languages. IWSLT 2024 was the first to host a shared task [^1] https://iwslt.org/2024/indic for AST of 3 Indic languages, but the findings showed that the narrow domain of the datasets made evaluation difficult Dabre and Song (2024).
Datasets for AST. The collection of speech-translation datasets for training end-to-end models has proven to be a persistent challenge. While efforts like SpeechMatrix Duquenne et al. (2022) and SeamlessAlign Communication et al. (2023) attempt to mine data from the web, the resulting datasets are often noisy. Most research has focused on curating data using comparable corpora from domains like Education Ďurišková et al. (2024); Song et al. (2019), TED talks Salesky et al. (2021), Parliament Debates Koehn (2005), among others. Additionally, with the rise of generative models, recent works have also started exploring the generation of large amounts of synthetic speech translation data in a flexible and cost-effective manner Bamfo Odoom et al. (2024). Although the use of synthetic data has been widely studied in text-based machine translation Sennrich et al. (2016); Gala et al. (2023), its application in Speech translation, particularly for low-resource languages like Indic languages, remains underexplored.
3 Indic-Spontaneous-Synth
lang Indic-En En-Indic
hours
utterances
hours
utterances
asm 3.17 1229 - -
ben 2.39 1137 9.28 2386
guj 0.73 302 9.29 4391
hin 2.04 901 9.32 2381
kan 2.08 815 9.14 4385
mal 2.30 995 9.22 4435
mar 1.39 609 9.30 4262
mni 1.31 772 - -
npi 3.08 1332 - -
ory 1.40 673 - -
pan 1.035 485 9.27 2545
tam 2.945 1695 9.22 4552
tel 2.634 852 9.13 4462
urd 1.786 796 9.28 2296
Total 28.29 12,593 92.45 36,095
Table 1: Statistics of data obtained across each language in both Indic-En and En-Indic directions for Indic-Spontaneous-Synth . For each language and in both directions, we show the number of hours and utterances.
Automatic Speech Translation (AST) systems are predominantly evaluated using the FLUERS benchmark Conneau et al. (2022), a speech-based version of the FLORES-101 Goyal et al. (2022) dev and dev-test sets, which spans 102 languages, including 14 Indian languages. This benchmark was created by recording human-translated English sentences from WikiSource, WikiTravel and WikiJunior, making it a representative of the Read speech type. However, this does not accurately reflect real-world day-to-day use cases where AST systems primarily encounter spontaneous speech  Fitzgerald et al. (2009).
Creating a benchmark for spontaneous, real-world speech translation is particularly challenging. Although there have been efforts to develop such datasets—such as Europarl-ST Iranzo-Sánchez et al. (2019), MuST-C Di Gangi et al. (2019b), and mTEDx Salesky et al. (2021)—there are no similar alternatives for Indian languages. To bridge this gap, we investigate the development of a synthetic benchmark utilizing LLMs, as illustrated in Figure 2. While our approach is not intended as a definitive solution, it emphasizes the current lack of resources for Indic languages rather than presenting an ideal benchmark. This method allows us to create a first-of-its-kind benchmark for Indian languages, following a similar approach used in previous synthetic benchmark efforts Lai et al. (2023). We use this as a stepping stone towards addressing the broader lack of resources in this area.
We introduce Indic-Spontaneous-Synth  a comprehensive benchmark covering 14 of the 22 scheduled Indian languages in the XX-to-English (XX-En) direction and 10 Indian languages in the English-to-XX (En-XX) direction. For the XX-En direction, we use IndicVoices Javed et al. (2024), a dataset that features natural, spontaneous speech and captures the cultural, linguistic, and demographic diversity of India. Since the test set of IndicVoices is not publicly available, we create Indic-Spontaneous-Synth using its dev set. We take the “normalized” transcripts from IndicVoices and prompt llama-3.1-405b-instruct model to generate their corresponding clean English translation. We deliberately chose not to use any specialized translation models like IndicTrans2 Gala et al. (2023), as these models are primarily trained for formal text-based translations and lack the capability to handle colloquial speech well. In contrast, with llama-3.1-405b-instruct we obtain steerable translations that perform well in informal and conversational contexts. Finally, we pair the English translations with their corresponding Indic audios, resulting in a dataset with a total of 42.53 hours of audio-text pairs across 14 Indian languages.
For the En-XX direction, we translate transcripts from the Svarah Javed et al. (2023) dataset, which is an Indian-accented English benchmark for ASR systems. We translate these transcripts to the top 10 Indian languages using llama-3.1-405b-instruct. We limit ourselves to the top 10 high-resourced Indic languages - Bengali (ben), Gujarati (guj), Hindi (hin), Kannada (kan), Malayalam (mal), Marathi (mar), Punjabi (pan), Tamil (tam), Telugu (tel) and Urdu (urd) - as we observed that llama-3.1-405b-instruct struggles to generate translations in lower-resource languages. However, it performs well in the XX-En direction due to its stronger English decoding capabilities Zhu et al. (2023); Alves et al. (2024). The translated utterances are then paired with the original English audio, resulting in 92.45 hours of English audio-Indic text pairs across 10 languages. Table 1 shows the overall statistics of Indic-Spontaneous-Synth.
While this synthetic approach has limitations, it serves as a crucial starting point in developing resources for spontaneous speech translation in Indian languages, a task that has not been extensively undertaken before. However, relying on existing models to generate and validate translations can introduce biases, which makes it difficult to fairly assess the performance of other models on such benchmarks. We chose llama-3.1-405b-instruct due to its large size and unsuitability for integration into AST systems, which at least helps mitigate bias in model evaluation. Nonetheless, we decided not to release the synthetically generated test set until it undergoes further human verification, which is currently in progress, to ensure the reliability of future evaluations.
4 Evaluating AST Systems
En-Indic Indic-En
lang SD SC S+IT2 W+IT2 SD W SC S+IT2 W+IT2 IW+IT2 IC+S IC+IT2
asm 37.9/—– 36.2/—– 40.5/—– 40.8/—– 44.0/19.6 18.8/9.4 46.3/20.7 47.1/25.0 18.3/6.3 —–/—– 45.3/16.2 47.2/33.6
ben 47.5/19.7 46.4/40.3 50.4/43.2 51.0/64.4 48.5/20.5 31.1/19.2 51.2/19.2 53.3/24.7 29.6/9.5 50.9/19.2 49.4/21.3 52.5/35.6
guj 48.9/20.5 48.4/39.8 51.7/41.9 52.1/62.8 52.8/27.6 35.3/20.8 55.9/28.8 57.5/31.2 34.1/11.9 54.3/21.2 55.9/28.9 57.1/35.0
hin 54.2/12.6 54.2/47.0 55.7/47.5 56.0/71.2 50.6/22.7 43.1/20.9 53.2/21.1 56.5/26.0 51.9/18.0 56.6/25.3 53.2/23.5 56.5/40.6
kan 48.9/21.5 47.8/41.2 51.4/43.2 52.1/65.0 47.6/14.8 28.6/11.2 50.1/13.6 52.0/16.4 37.5/5.5 51.4/9.8 50.0/10.9 52.0/29.1
mal 49.0/20.1 47.9/35.5 53.9/41.0 54.7/60.5 48.4/18.4 33.8/15.4 52.0/18.5 54.1/21.4 18.9/9.9 52.8/13.3 52.2/19.4 54.2/33.9
mar 44.1/15.0 42.3/35.3 47.2/39.6 47.5/59.5 48.4/22.5 30.6/19.3 52.2/22.6 53.5/27.6 39.5/11.9 52.8/23.4 52.5/25.5 54.3/34.5
npi 50.0/—– 47.2/—– 53.1/—– 53.7/—– 50.6/17.8 32.2/14.7 54.3/18.4 56.4/20.0 40.4/11.4 —–/—– 52.8/19.0 55.3/29.5
ory 46.2/—– 43.4/—– 46.5/—– 46.8/—– 48.6/30.1 24.9/19.6 52.8/29.2 54.2/32.7 17.4/8.5 49.6/20.9 53.5/30.7 55.1/35.2
pan 50.0/26.5 48.8/40.7 50.9/43.1 51.5/65.1 50.3/23.3 32.8/24.5 53.0/26.9 54.7/29.5 26.2/10.7 51.4/21.3 54.1/27.2 55.1/41.3
snd 47.1/—– 45.9/—– 42.0/—– 42.3/—– 31.1/10.9 24.5/12.8 36.2/9.5 38.5/5.5 17.6/5.6 —–/—– 18.2/10.1 39.3/23.5
tam 50.6/16.0 49.5/38.4 53.0/42.2 53.3/63.5 44.8/13.6 27.5/14.5 48.1/11.8 50.3/13.2 42.3/10.6 49.9/15.8 47.2/14.2 50.7/24.4
tel 52.0/18.5 51.3/39.7 54.5/43.2 55.3/64.7 47.5/19.4 31.7/12.8 51.5/16.7 54.4/20.3 22.1/6.2 54.1/13.0 51.7/15.4 55.6/36.6
urd 47.2/25.6 46.6/42.4 48.8/44.4 48.8/67.5 47.6/24.8 39.0/24.5 50.3/25.9 53.1/34.6 49.0/26.8 48.7/27.2 49.5/24.4 51.3/48.4
avg 48.1/19.6 46.9/40.02 49.9/42.9 50.4/64.4 47.2/18.9 30.9/14.8 50.5/18.6 52.5/21.5 31.7/10.9 52.0/19.1 48.9/20.4 52.5/31.4
Table 2: ChrF++ scores of all the AST systems evaluated on the FLUERS and Indic-Spontaneous-Synth set in the En-Indic and Indic-En directions. Each entry in the table represents two ChrF++ scores in the format $X/Y$, where $X$ corresponds to the score on FLUERS and $Y$ corresponds to the score on Indic-Spontaneous-Synth. The best-performing systems are bolded for each direction, each language and each benchmark. Higher value shows higher performance.
In this section, we present our findings after evaluating various speech translation systems on Indic Languages. We begin by describing the experimental setup - description of various systems under the Direct and Cascaded paradigms - followed by the discussion of the results.
4.1 Experimental Setup
In this study, we primarily evaluate the Unified (or end-to-end) and the Cascaded Speech translation paradigms across various models. While recent advancements in Audio-LLMs (or SLLMs) introduce new paradigms by integrating Speech Foundation Models with LLMs for Automatic Speech Translation (AST), we do not include them in this study due to the lack of support for Indian languages.
Unified Systems:  In this paradigm, a single end-to-end speech translation model is used. The unified AST models considered in our study are as follows:
SeamlessM4T (SD) [Uncaptioned image] Communication et al. (2023) is a foundational multilingual and multitask model supporting speech-to-speech (S2S), speech-to-text (S2T), text-to-speech (T2S), text-to-text (T2T) translation, and automatic speech recognition (ASR) for up to 100 languages. For our experiments, we use the Seamless-M4T-v2-large model.
Whisper (W) [Uncaptioned image]Radford et al. (2023) is a multilingual model trained for automatic speech recognition (ASR) and speech translation (AST). For our experiments, we use the Whisper-large-v3 model which has been trained on 1 million hours of weakly labeled and 4 million hours of pseudo-labeled audio.
Cascaded Systems:  In this paradigm, separately trained Automatic Speech Recognition (ASR) and Machine Translation (MT) models are combined in a pipeline for performing Automatic Speech Translation (AST). The different ASR and NMT models used in our experiments are listed below.
SeamlessM4T(SC)[Uncaptioned image]Communication et al. (2023) where we use the ASR and NMT capabilities of the SeamlessM4T model [^2] https://github.com/facebookresearch/seamless_communication .
Whisper (W) [Uncaptioned image]Radford et al. (2023) where we use the ASR capability of the Whisper model.
Indic-Whisper (IW)[Uncaptioned image]Bhogale et al. (2023) is a fine-tuned Whisper model [^3] https://github.com/AI4Bharat/vistaar supporting only 12 Indian Languages.
Indic-Conformer (IC) [Uncaptioned image] which is a suite of ASR models [^4] https://github.com/AI4Bharat/IndicConformerASR based on the Conformer architecture, models supporting all 22 scheduled Indic Languages.
IndicTrans2 (IT2) [Uncaptioned image]Gala et al. (2023) which is an open-source transformer-based multilingual NMT model supporting translations across all 22 scheduled Indic languages. For our experiments, we use the IndicTrans2-1B model [^5] https://github.com/AI4Bharat/IndicTrans2.
In our experiments, we combine the IndicTrans2-1B model and the Seamless-M4T V2 Large translation model with various ASR models to perform automatic speech translation (AST). Specifically, the translation capabilities of IndicTrans2-1B and Seamless-M4T V2 Large (T2TT) are integrated with ASR outputs from systems such as Whisper, Indic-Whisper, and Indic-Conformer.
4.2 Metrics
Various metrics have been proposed over the years to assess the quality of translation systems, such as Word Error Rate (WER), BLEU Papineni et al. (2002), COMET Rei et al. (2020) among others. chrF++ (Popović, 2017), an extension of the chrF metric (Popović, 2015) that additionally considers word unigrams and bigrams, and is better correlated with human judgments. In all our experiments, we report the chrF++ scores using the SacreBLEU Post (2018) package due to its ability to handle the rich morphological variations of Indic Languages. For Indic-En [^6] Indic-En sacreBLEU ChrF++
signature:nrefs:1|case:mixed|eff:yes|nc:6|nw:2|space:no|version:2.3.1
evaluation, we use the in-built default mteval-v13a tokenizer, while for En-Indic [^7] En-Indic sacreBLEU ChrF++
signature:nrefs:1|case:mixed|eff:yes|nc:6|nw:2|space:no|version:2.3.1
evaluation, we use Indic tokenizers from IndicNLP Kunchukuttan (2020) and Urduhack [^8] https://github.com/urduhack/urduhack libraries to tokenize the machine translations and reference translations before running SacreBLEU Post (2018).
4.3 Results and Discussion
We report the performance of various Automatic Speech Translation (AST) systems evaluated on the Fleurs benchmark and on Indic-Spontaneous-Synth benchmark, as summarized in Table 2.
Effectiveness of Cascaded Models: A key trend across experiments is the strong performance of language-specialized models in both ASR and MT tasks. In the En-Indic direction, the Whisper ASR model cascaded with IndicTrans2-1B performs particularly well, while in the Indic-En direction, the Indic-Conformer model combined with IndicTrans2-1B achieves the best results. This strong performance can be attributed to the models’ enhanced ability to handle disfluencies and the variability typical of spontaneous speech.
Performance Challenges in Translating Spontaneous vs. Read Speech: Another consistent trend is the significant drop in performance when translating spontaneous speech in Indic-Spontaneous-Synth as compared to the read speech in Fleurs, particularly in the Indic-En direction. Although cascading specialized models helps mitigate this issue to some extent, the performance gap remains large, especially in direct models. These results indicate that while current models perform reasonably well on read speech, they struggle with spontaneous speech, underscoring the need for improved datasets, training methodologies, and more robust evaluation frameworks that address the challenges of real-world speech.
Difficulties in processing local accents: Another key limitation observed is the poor performance of some models, notably SeamlessM4T, on Indian-accented English in Indic-Spontaneous-Synth. Upon examining the outputs, we found that the model often transcribes the input instead of translating it, suggesting a lack of Indian-accented data in its training corpus.
These findings highlight the importance of developing a diverse, comprehensive dataset that covers spontaneous speech across various domains and tasks. Such a resource would enable systems to be more effectively trained and evaluated for real-world speech scenarios.
5 BhasaAnuvaad
lang Existing Mined Synthetic
Indic-TedST VaaniPedia WordProject UGCE Resources Mann Ki Baat NPTEL Indic Voices
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
asm - - 3.4 1.6 - - - - 28.3 15.7 10.6 5.3 348.9 23.5
ben 13.1 6.5 121.0 52.5 50.7 23.9 431.7 329.6 38.3 21.4 1516.6 618.8 300.5 30.3
guj 17.2 1.5 121.0 52.5 50.7 23.9 431.7 329.4 39.1 22.0 1977.9 784.9 16.7 2.8
hin 100.4 52.5 121.0 52.5 50.7 23.9 431.7 329.2 37.7 21.3 3021.6 1176.8 206.6 20.4
kan 3.7 1.9 23.2 9.9 50.7 23.9 431.7 328.6 36.9 20.8 1601.4 638.0 175.6 11.7
mal 9.2 3.2 0.5 52.5 50.7 23.9 431.7 329.6 38.3 21.3 1729.1 699.3 278.1 22.9
mar 29.0 12.7 56.5 24.4 50.7 23.9 431.7 322.4 39.0 22.0 1601.7 626.3 138.0 16.2
mni - - - - - - - - 29.7 16.7 - - - -
npi - - 121.0 52.5 - - - - - - - - 343.9 31.6
ory - - 79.0 34.0 49.2 23.1 - - 36.8 20.5 0.8 0.3 161.1 19.1
pan 1.2 0.4 0.5 0.3 50.7 23.9 - - 37.0 20.5 9.6 5.8 194.1 18.9
tam 20.1 10.1 121.0 52.5 30.5 14.1 431.7 327.9 37.4 20.9 2631.9 1048.1 378.6 33.8
tel 9.1 2.2 121.0 52.5 50.7 23.9 431.7 322.6 37.8 21.4 1839.9 741.2 314.3 21.6
urd - - 2.3 1.1 50.7 23.9 - - 29.0 16.0 - - 186.6 19.1
Total 202.9 91.0 891.1 439.0 536.2 252.3 3453.8 2619.3 465.4 260.7 15940.9 6344.7 2694.1 269.0
Table 3: Overview of En-Indic spoken translation dataset across various Indic languages. The table details the hours of audio and the number of utterances collected from multiple sources. Number of utterances are in thousands (K)
lang Existing Mined Synthetic
CVSS
& Khan
Academy
Seamless Data WordProject Spokentutorial Mann Ki Baat NPTEL Indic Voices
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
hours
uttr.
asm - - - - - - 36.9 23.2 23.0 13.9 9.3 5.7 443.6 26.0
ben - - - - 117.6 28.9 49.9 31.1 38.5 20.7 388.4 149.1 450.6 33.2
guj 12.0 - - - 82.1 29.0 63.6 37.6 37.4 21.3 795.0 300.9 30.2 3.0
hin - - 2355.4 1021.5 88.5 29.5 63.6 37.9 40.9 21.3 1433.9 477.5 282.9 21.6
kan - 174.2 68.6 108.7 29.5 25.0 15.4 38.5 21.1 569.9 219.3 199.5 12.2
mal - - - - 134.4 55.5 24.4 15.4 35.6 24.1 721.0 320.2 378.4 24.9
mar - - - 98.8 29.2 68.9 41.3 42.1 22.7 738.7 289.7 201.7 17.5
npi - - - - - - 58.8 35.4 - - - - 479.1 34.0
ory - - - - 97.1 28.5 6.5 3.5 34.0 20.0 - - 219.6 20.0
pan - - - - 113.2 1.1 19.8 11.4 78.2 8.2 - - 231.4 19.9
tam 96.0 2.5 1222.4 478.8 51.5 18.9 69.1 39.9 41.2 19.9 1231.8 538.8 454.5 35.5
tel - - 873.3 329.5 90.4 29.2 12.5 6.5 35.3 26.3 738.4 302.1 377.5 22.8
urd - - 2837.1 1107.7 94.3 29.4 - - 11.9 7.0 - - 239.2 17.4
Total 108.0 62.8 7462.3 3006.0 1076.7 308.6 499.0 298.7 456.4 226.5 6626.5 2603.3 3988.2 288.0
Table 4: Overview of Indic-En spoken translation dataset across various Indic languages. The table details the hours of audio and the number of utterances collected from multiple sources. Number of utterances are in thousands (K)
To overcome the challenges with spontaneous speech, we introduce BhasaAnuvaad , a large-scale, high-quality dataset covering 14 of the 22 scheduled Indian languages across a wide array of domains, use cases and styles.
BhasaAnuvaad  is the largest spoken translation dataset for any Indian language, totaling approximately 44,400 hours of speech and text data. Its construction involved a multi-step approach, as illustrated in Figure 2, combining:
  1. 1.
    Aggregating existing datasets
  2. 2.
    Mining parallel speech and text data from comparable sources, and
  3. 3.
    Synthetically generated speech/text parallel data
A language wise composition of this dataset is illustrated in Figure 3. Tables 3 and 4 detail the composition of data from various sources across the 10 languages within these 3 categories.
5.1 Aggregating Existing Datasets
We began by aggregating several existing datasets in Indian Languages, which form the foundation of BhasaAnuvaad . These include:
Indic-TEDST: The Indic-TEDST Sethiya et al. (2024) dataset includes TED Talks audio, transcripts, and translations in nine widely spoken Indian languages—Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Punjabi, Tamil, and Telugu—totaling approximately 203 hours of English-to-Indic data.
CVSS: The CVSS Jia et al. (2022) corpus is a multilingual speech-to-speech translation dataset derived from CoVoST 2, which itself originates from the Common Voice dataset. Spanning 21 languages, it includes only one Indian language, Tamil, with approximately 3 hours of Tamil-English data.
Khan Academy Corpus: The Khan Academy corpus Ďurišková et al. (2024) provides Indic-En data in 29 languages. It includes 3 Indian languages, offering over 300 hours of audio, primarily in the educational domain.
SeamlessAlign: We recreate the SeamlessAlign data introduced in Communication et al. (2023) following the same pipeline to get around 7500 hours of data for 5 Indic languages.
5.2 Mining from comparable sources
We explore the mining for parallel speech-text data from various comparable online sources that may contain high-quality parallel pairs. These sources include publicly available audio and video content spanning diverse domains, with a focus on religious and educational materials. Although many sources provided transcripts in different languages, we found that the transcripts were often not properly aligned with the corresponding audio. To address this issue, we developed a pipeline to align the transcripts and audio, creating high-quality parallel speech-text data.
5.2.1 Cleaning and Alignment Pipeline
Our cleaning and alignment pipeline processes the audio and texts in four stages: Normalization, Punctuation Restoration, Alignment, and Sentence Pair Extraction.
  1. 1.
    Normalization: Audio files are first converted to mono channel and resampled to 16,000 Hz. The transcript files are also formatted and cleaned by removing line breaks, extra spaces, and other unnecessary symbols. Terminal punctuations are replaced with a fixed non-standard character to enable easier sentence splitting during alignment.
  2. 2.
    Punctuation Restoration: Since subtitles and transcripts often omit punctuation, we use the DeepMultilingualPunctuation model [^9] https://github.com/oliverguhr/deepmultilingualpunctuation, to restore relevant punctuations and produce cleaner text.
  3. 3.
    Audio-Transcript Alignment: The processed data is then passed to the Nemo Forced Aligner (NFA)  [^10] https://docs.nvidia.com/nemo-framework/user-guide/latest/nemotoolkit/tools/nemo_forced_aligner.html which aligns transcripts with their corresponding audio. NFA generates token-, word-, and segment-level timestamps using CTC-based ASR models. For Indic audios, we use the IndicConformer models, while for English, we utilize the Nvidia FastConformer [^11] https://huggingface.co/nvidia/stt_en_fastconformer_ctc_large. The non-standard character introduced during normalization is used to extract sentences from transcripts. To handle longer audio files effectively, we modify NFA to first divide the audio into chunks using the Silero VAD Team (2024). Each chunk is processed separately, and the resulting logits are merged for final alignment. Alignment quality is evaluated using Levenshtein distance similarity scores between the reference and predicted sentences, following the approach by Bhogale et al. (2022).
  4. 4.
    Parallel Text Alignment: At this stage, we have audio segments paired with their corresponding transcripts. For aligning these with corresponding target language texts (Indic languages for the English-Indic direction and vice versa), we split the target transcripts into sentences and obtain the multilingual embeddings of both the source audio transcripts and the target sentence transcripts using SONAR Duquenne et al. (2023). We then calculate the cosine similarity scores between each pair of embeddings to determine the parallel pairs. The target language transcripts with the highest cosine similarity to each source audio segment are identified as parallel pairs. We do not apply any filtering based on fixed thresholds; instead, we release the alignment and cosine similarity scores so users can filter the data according to their own criteria.
Figure 3: Distribution of dataset hours in log scale for En-Indic (left) and Indic-En (right) language pairs across three data sources: Previous Datasets (blue), Curated & Mined (green), and Synthetic (red).
5.2.2 Sources of Mining Corpora
Spoken Tutorial: Spoken tutorial [^12] https://spoken-tutorial.org/, is an initiative under the National Mission on Education under the Ministry of Human Resource Development, Government of India to enable the translation of various vocational educational content in different languages. This dataset covers Indic-En translations with over 499.04 hours of educational tutorials in 12 Indian languages.
UGCE Resources: The University Grants Commision [^13] https://ugceresources.in/ translates various University level courses to different languages. The courses are taught in English (En audio) and the transcripts are translated to various 8 other languages, contributing over 430 hours of audio data.
VaaniPedia: Vaanipedia [^14] https://vanipedia.org/ is a large collection of religious and spiritual, En audio-Indic text content in 15 Indian languages covering over 122 hours of audio.
Mann ki Baat: Mann Ki Baat [^15] https://www.pmindia.gov.in/ is a weekly radio podcast organized by the Indian Prime Minister to discuss various national Issues. The podcast is transcribed and manually translated into 13 Indian languages. It is then recorded in these languages, also through manual efforts, resulting in over 425 hours of speech data for both Indic-En and En-Indic directions. Notably, Mann Ki Baat is the only source in this dataset that includes Manipuri, though it is presented in Bengali script.
WordProject: This dataset consists of Bible audiobooks sourced from the Word Project website [^16] https://www.wordproject.org/ in 12 Indian languages, totaling approximately 1,200 hours of speech data. It includes content in both Indic-En and En-Indic directions.
NPTEL: The National Programme on Technology Enhanced Learning (NPTEL) [^17] https://nptel.ac.in/ is an Indian e-learning platform for university-level science, technology, engineering, and mathematics (STEM) subjects. It is the largest e-repository in the world, offering courses in engineering, basic sciences, and selected humanities and management subjects. These courses are primarily recorded in English, but they are transcribed and re-recorded in multiple languages manually. This is the largest component of our dataset, contributing over 22,500 hours of spoken content across 10 Indian languages. It covers a wide range of STEM subjects and includes both En-Indic and Indic-En translations, making it one of the most significant sources in our collection.
While these datasets are high quality, they are not diverse with majority of our content being from Educational domain. Furthermore, the audio styles they cover lack the diversity necessary for models to adapt effectively to real-world, where background noise, spontaneous speech, and varying accents are common. To overcome these limitations, we augment our dataset by using synthetic data generation.
5.3 Synthetic Data Generation
To ensure that we have audio diversity across different domains, demographics, and languages, we use IndicVoices which is the largest collection of natural and spontaneous speech containing over 4000 hours of transcribed speech. Its extensive language diversity makes it well-suited for training models capable of performing effectively in real-world scenarios.
Synthetic Translations: Similar to Indic-Spontaneous-Synth we use the llama-3.1-405b-instruct model to translate the Normalized transcripts from IndicVoices  train set following the process outlined in Section 3. We then pair the Indic audio with the corresponding transcripts to get the Indic audio-English text pairs.
Synthetic Speech Generation: We also explore the use of synthetically generated speech, as manual recording with voice artists can be time-consuming and costly. We take the verbatim transcripts of IndicVoices and prompt the llama-3.1-405b-instruct model to generate their colloquial translations that reflect spontaneous speech, including filler words and natural hesitations. We then use an internal Parler-TTS Lacombe et al. (2024); Lyth and King (2024) model fine-tuned on Indian languages to generate English audios. Since this model has been fine-tuned on Indian audio, we get high-quality Indian-accented English audio. Finally, the generated English audio is paired with the normalized transcripts from IndicVoices to form English audio-Indic text pairs.
Building on the insights gained from evaluating existing systems and recognizing their shortcomings, we introduced BhasaAnuvaad —a comprehensive, large-scale dataset designed to bridge the performance gap in real-world spoken translation applications. With BhasaAnuvaad, we aim to address the complexities of spontaneous speech across the 14 of the 22 scheduled languages. By incorporating diverse domains and varied speaking styles, BhasaAnuvaad  aims to support the development of models capable of meeting the dynamic requirements of deployment environments. Through this resource, we aim to not only enhance the robustness of Spoken Translation systems but also provide a foundational tool that supports future advancements in multilingual speech translation for Indian languages.
6 Conclusion
In this work, we evaluated popular Spoken Translation systems for Indian languages and identified key limitations in their ability to handle spontaneous speech, a common occurrence in real-world deployment scenarios. Our findings underscore the need for more rigorous evaluation and benchmarking of these systems in practical settings. To address this gap, we introduced a novel benchmark,  Indic-Spontaneous-Synth, comprising 130 hours of spontaneous speech data across 10 languages, covering diverse domains and tasks to provide a more realistic testbed for system performance. Additionally, recognizing the scarcity of resources covering multiple languages, domains, and speaking styles, we presented  BhasaAnuvaad-—a large-scale, meticulously curated dataset featuring over 44,000 hours of speech data and 17 million text segments across 10 languages. This dataset includes a broad spectrum of practically relevant domains and styles, offering a robust foundation for training models.
This paper is a work in progress. Future work will involve expanding this dataset to include additional low-resource Indian languages and exploring the effectiveness of both cascaded and direct models on this data to further improve translation accuracy. We believe these contributions will significantly enhance the robustness and applicability of spoken translation systems for Indian languages, paving the way for more accurate, scalable solutions that address diverse linguistic and contextual needs.
Acknowledgements
We would like to thank the Centre for Development of Advanced Computing (CDAC [^18] https://www.cdac.in/index.aspx?id=pune ), Pune for generously supporting this work and providing us access to multiple GPU nodes on the Param Siddhi Supercomputer. We would like to thank the EkStep Foundation and Nilekani Philanthropies for their generous grant which went into hiring human resources as well as cloud resources needed for this work. We would also like to thank Srihari Krishna of AI4Bharat for creating the BhasaAnuvaad logo.